OpenStudy (anonymous):

What are the vertices of the hyperbola given by the equation y^2 – 9x^2 + 22y – 72x – 32 = 0? (-3, -11) and (-5, -11) (-1, -11) and (-7, -11) (-4, -8) and (-4, -14) (-4, -10) and (-4, -12)

OpenStudy (helder_edwin):

complete the squares first!

OpenStudy (anonymous):

y^2+81x+22y-72y-32=0?

OpenStudy (helder_edwin):

let's see. u have \[ \large y^2-9x^2+22y-72x-32=0 \] \[ \large (y^2+22y+\qquad)-9(x^2+8x+\qquad)=32 \]

OpenStudy (helder_edwin):

what numbers would u put in the blanks???

OpenStudy (helder_edwin):

@Bunnijjane ???

OpenStudy (anonymous):

the second blank is 5.65 i think. I was trying to figure out the first.

OpenStudy (helder_edwin):

no

OpenStudy (helder_edwin):

for the first it would be \[ \large \left(\frac{22}{2}\right)^2=? \]

OpenStudy (anonymous):

121?

OpenStudy (helder_edwin):

yes. now do the same for the second part!!

OpenStudy (anonymous):

16?

OpenStudy (helder_edwin):

great!!

OpenStudy (helder_edwin):

now we have \[ \large (y^2+22y+11^2)-9(x^2+8x+4^2)=32+\qquad-\qquad. \]

OpenStudy (helder_edwin):

u have to put two numbers in the blanks. which are they?

OpenStudy (helder_edwin):

the ones u just added!!

OpenStudy (anonymous):

wow I'm a moron -_-. Sorry! so now the equation is (y^2+22y+11)-)x^2+8x+4^2)=32+121-12?

OpenStudy (helder_edwin):

the first is correct.

OpenStudy (helder_edwin):

the second one would be \[ \large -9(16)=-144 \]

OpenStudy (helder_edwin):

do u get this??

OpenStudy (helder_edwin):

so the equation would be \[ \large (y+11)^2-9(x+4)^4=32+121-144=9 \] dividing by 9 we have \[ \large \frac{(y+11)^2}{9}-\frac{(x+4)^2}{1}=1 \]

OpenStudy (anonymous):

The answer is c.(-4, -8) and (-4, -14)